# Continual Learning in Medical Imaging from Theory to Practice: A Survey and Practical Analysis

### v1

@@ 3.1.2 Functional Regularization
Kim et al. (2018) proposed learning an inverse function that maps the model’s logits to the feature space back and using it to regularize the feature space of the model in the next task. They demonstrated the effectiveness of their method in a multidomain, single-task X-ray classification scenario. An interesting domain incremental study provided a dual-distillation method combined with a replay to mitigate forgetting (Li et al., 2020). Dual-distillation is carried out using the logits from the old and a new fully fine-tuned model to train an updated model that performs well on new and old tasks. Recent work on manifold learning Akundi and Sivaswamy (2021) used a reformer network to augment any input domain image to lie in the same data manifold. This, combined with various regularization-based methods, was shown to be effective in reducing forgetting due to domain shifts in the clinical data. Moreover, a multimodal study by Patra et al. (2021) on ultrasound and eye-tracking data utilized distillation to train a model on a continual stream of data, whereas Patra et al. (2021) proposed an architecture that leverages the simultaneous availability of two or more datasets to learn a disentanglement between the content and domain in an adversarial manner. This allows a domain-invariant content representation that lays the basis for the continual semantic segmentation of MRI images.He et al. (2021) use a probability-map alignment scheme to integrate the previous and current map for a segmentation task. They further used knowledge distillation to transfer the knowledge from the integrated probability map to the current model, thereby improving the segmentation performance of the current model in the CL. Chakraborti et al. (2021) presents a semi-supervised continual learning method combined with a contrastive loss. Their approach was to first train an auto-encoder with a small amount of labeled data in the usual supervised manner, followed by an unsupervised adaptive regularization-based contrastive learning approach.
@@
Another work by Liu et al. (2022b) introduced a method for multi-organ segmentation in CT images. Their methodology uses lightweight memory modules and a tailored regularization loss function, allowing the model to effectively utilize replayed examples to prevent forgetting. Here, the introduced loss function constrains the learned feature space to maintain the feature of an old class close to its mean representation within the memory module, while ensuring that the features of a new class are distinct from all representations of the old class. Furthermore, Tian et al. (2022) proposed a multi-scale multitask CL framework. Their method combined contrastive learning with distillation-based regularization and data rehearsal to enhance learning without catastrophic forgetting. Li et al. (2022a) used feature whitening to regularize the important domain in-variant features in a domain incremental cardiac image segmentation study. This feature regularization was augmented by a generative feature reply, where they proposed a base conditional generator along with a style module to produce images for replay. Furthermore, Roy et al. (2023) introduced the use of a mixed-curvature embedding space for knowledge distillation to prevent catastrophic forgetting. The authors employed a combination of Euclidean and hyperbolic spaces to project the learned feature embeddings, emphasizing their importance for complex data, especially in medical imaging. Another recent study by Gao et al. (2023) used a Kullback–Leibler (KL) divergence-based loss function to preserve the models’ knowledge by penalizing the change in output of the new and old models. More recently, the Fourier test–time adaptation (FTTA) method was presented for classifying medical images with domain gaps (Huang et al., 2023). This technique utilizes a Fourier-based domain adjustment strategy along with a multilevel consistency approach to regularize the model’s output. The proposed method is based on a multilevel consistency measurement designed for self-correcting predictions.
Overall, regularization-based methods offer an efficient and privately protected approach to address catastrophic forgetting in CL. These methods do not introduce any computational overhead or privacy concerns, making them well suited for real-life medical applications. However, standalone regularization-based CL methods often fall short of incrementally learning very different tasks, such as diverse class incremental learning, whereas their effectiveness becomes more pronounced in domain incremental settings. Recent attempts to combine regularization-based CL methods with other CL techniques have been beneficial in addressing this issue, which can be a promising direction for making regularization-based CL methods practical for medical applications.
3.2 Replay-Based Approaches
@@
Venkataramani et al. (2019) use Memory-augmented neural networks(MANN) for continuous domain adaptation for the task of semantic segmentation. They proposed a single framework for domain adaptation in which images similar to the target image and their respective textual and shape features are fetched using wavelet features of the target image as a query. These context features, along with the target image features, were embedded into the model trained using only the source domain. Ravishankar et al. (2019) proposed a single approach for the different variants of lifelong learning, such as multi-task setting and single incremental task setting, aiming to ensure data privacy and mitigate catastrophic forgetting. They employed a pseudo-rehearsal strategy using a finite memory module and ensured class separation using a composite loss. Furthermore, they introduce feature transformers to learn new representations from the old data as well as the new data. Hofmanninger et al. (2020a) proposed a rehearsal method that infers the data shifts by adopting a dynamic memory module. Their memory update strategy was designed based on a fixed set of rules, and the features were replaced with a high-level metric calculated using gram matrices. Perkonigg et al. (2021) employs the same strategy but also introduces an additional module called the pseudo-domain (PD) model, which detects clusters of similar styles from the continuous stream of data. These pseudo-domains are seen as proxies for the unknown, real domains, and are used to balance the memory and training process (DM-PD). Srivastava et al. (2021) also proposed another domain incremental learning that leverages vector quantization approach under limited memory constraints to effectively store and replay hidden representations where no clear domain-shift boundaries exist and no a-priori knowledge is possible. Wang et al. (2021) introduced a method that integrates medical expertise into disease diagnosis by considering context and medical entity characteristics in order to transfer knowledge to new stages. Their work was among the pioneering studies exploring CL within medical domains, specifically focusing on diagnosing diseases based on clinical notes. Additionally, they introduced a new technique called embedding episodic memory and consolidation (E2MC) to avoid catastrophic forgetting in disease diagnosis tasks. Morgado et al. (2021) delve into continual learning within dermatology, employing four distinct methods: naive, elastic weight consolidation, averaged gradient episodic memory, and experience replay. These methodologies allow adaptation to new conditions while safeguarding previously acquired knowledge in the field.
@@
To overcome the problem of insufficient sensitive patient data for digital biomarker model training owing to privacy laws, Shevchyk et al. (2022) proposed a Privacy-Preserving Synthetic Data Generation pipeline that uses a GAN to generate anonymous medical samples. Privacy is ensured using a Siamese Neural Network (SNN) with triplet loss to evaluate the generated data. The effectiveness of this method was demonstrated using a Respiratory Sound Classifier that collects respiratory symptom information and is relevant for conditions such as Asthma, COPD, or COVID-19.
@@
Similarly, Liu et al. (2022a) pioneered a generative replay-based approach in semantic segmentation tasks for medical imaging by introducing SegGAN to prevent catastrophic forgetting. This GAN generates both previous images and their corresponding pixel-level labels, addressing privacy limitations while preserving the acquired knowledge for the subsequent learning stages. Additionally, they proposed a unique filtering mechanism that ensures the selection of high-quality generated data by maintaining consistency between the pseudo-labeling and generative replay techniques. Li et al. (2022c) propose the use of multiple generators to improve the quality of replayed samples. Instead of using generators to generate higher-resolution images, this study uses a fixed feature extractor and focuses on feature generation, which requires fewer parameters and fewer computations. Li et al. (2022b) introduced a fresh approach for CL in gesture recognition, specifically for adding new gesture classes. This method focuses on managing out-of-distribution (OOD) memory and improving rehearsal to consistently learn new gestures. Additionally, it incorporates an energy-based loss to penalize out-of-distribution gesture examples and further refine the model during training. Ayromlou et al. (2022) introduces an innovative data-free framework for incremental class learning. This approach initially creates synthetic data from the model trained on past classes to form a Class Impression, and then updates the model by merging the synthesized data with new class data. They employ various loss functions, such as cosine normalized cross-entropy, margin loss, and intra-domain contrastive loss, to tackle class imbalances, differentiate between previous and new classes, and ensure the model’s generalization from synthetic to real data.Zhang et al. (2022) introduce a new method, SMG (Synchronous Memorizability and Generalizability) -learning, to enhance both memorization and generalizability. It presents the Synchronous Gradient Alignment (SGA) objective of improving network memorization and adaptability across previous and new sites. To streamline optimization without excessive computational load, the dual-meta algorithm was developed. Moreover, to enhance rehearsal effectiveness, the replay buffer is modified to include diverse data from various sites, thus minimizing redundancy.
@@
Bera et al. (2023) introduce a novel solution to improve image selection to create the memory bank by ranking and selecting images based on their contribution to the learning process. In situations such as medical applications where storing previous data is limited by privacy regulations, Chen et al. (2023) presented Generative Appearance Replay for continual Domain Adaptation (GarDA), an approach employing generative replay techniques. GarDA facilitates the gradual adjustment of a segmentation model to new domains by utilizing unlabeled data. Similarly, Thandiackal et al. (2023) proposed a generative feature-driven image replay with a dual-purpose discriminator to align features for continual learning in unlabeled domains, achieving state-of-the-art results in histopathology datasets. Unlike single-step unsupervised domain adaptation (UDA), continual adaptation across various domains enables the utilization and consolidation of knowledge from different domains throughout consecutive learning phases. Wei et al. (2023) proposes a data selection strategy using Variational autoencoders (VAE) and adversial network. Through this approach, they retain training efficiency by always selecting a fixed subset size from the entire dataset. Kim et al. (2024) proposes a continual learning for multicenter studies, eliminating the need for a central server and mitigating the risk of catastrophic forgetting of previously acquired knowledge. The framework involves a method selection process for continual learning, acknowledging that no single method suits all datasets and utilizes fake data generated by a generative adversarial network to evaluate the methods.
Replay-based techniques can be used in medical applications; however, doing so poses additional challenges with regard to model interpretability, regulatory framework validation, and ethical issues that must be addressed to ensure the safety and effectiveness of AI-driven healthcare solutions. Addressing these challenges is crucial for harnessing the full potential of CL techniques in the medical domain while ensuring patient safety and privacy.
3.3 Dynamic Model Approaches
@@
Previous methods were either limited by the model’s parameter change in subsequent tasks or relied on rehearsal techniques. A more recent and promising solution involves intelligent introduction of new parameters for each task (Figure 6. In other words, shared features are maintained among tasks to avoid repeated training, similar to ensembling. Moreover, employing a set of new parameters enables more effective learning of task-specific features.
@@
Dynamic models offer a direct solution by allocating specific segments of the model to distinct tasks. A study by Karani et al. (2018) addressed domain shifts in lifelong learning scenarios using domain-specific batch normalization layers for MRI segmentation. Variance in MRI image intensities, caused by multiple factors, leads to this domain shift (Preboske et al., 2006; Jovicich et al., 2006).
@@
However, another challenge arises when we lack knowledge about the task or which sub-network to choose during inference for particular data. For example, if a model is trained on images from three different scanners, determining the domain during inference requires image metadata to activate the corresponding model segment. Gonzalez et al. (2020) tackled this issue by training separate UNet models for domain segmentation. To discern the domain during inference, they utilized separate Variational Autoencoders (VAEs) for each domain and selected the domain with the minimal VAE reconstruction loss. Another study by González et al. (2022) introduces a task-agnostic continual learning approach applicable across diverse deep learning architectures. Their method involved utilizing the Mahalanobis distance to identify domain changes. Upon detecting such changes, they expand the network, ensuring a nonlinear expansion. During inference, the network leverages this distance measurement to determine task ID, thereby facilitating efficient task identification. Similarly, Bayasi et al. (2023) selected a network based on a distance-based metric for each test dataset. In contrast, Zhao et al. (2023) used different prompts for each task and used a matching strategy to choose the correct prompt at inference.
@@
Recently, advanced techniques have been applied to address the challenges of multiorgan segmentation. These methods operate in a class incremental setting, requiring the model to segment all the regions observed in previous tasks during inference. Zhang et al. (2023c) highlights the significance of dynamically expanding models to encompass new data and classes for segmenting multiple organs and tumors. Their architecture features lightweight class-specific heads that replace traditional output layers, enabling separate predictions for both new and previously learned classes. Leveraging extensive image-text co-training, the authors embedded contrastive language-image pretraining (CLIP) embeddings into organ-specific heads to capture each class’s semantic essence. Research by You et al. (2022) introduced an innovative incremental transfer learning framework designed to progressively address multi-site medical image segmentation tasks. They proposed strategies at both the model and site levels aimed at incremental training to enhance segmentation, improve generalization, and boost transfer performance, particularly in settings constrained by limited clinical resources.
@@
Ji et al. (2023) introduced an innovative architectural framework for semantic segmentation. Their approach learns a unified deep segmentation model capable of segmenting 143 whole-body organs by training a single unified deep segmentation model. For each segmentation task, a trained encoder remained fixed, whereas a new decoder was gradually added. To handle model complexity, they utilized a progressive pruning strategy involving neural architecture search and teacher-student knowledge distillation. This strategy helps maintain performance while controlling the size of the model.
@@
The practice of dynamically adjusting models extends to classification tasks, as observed in a recent study by Zhang et al. (2023b) on various medical imaging classification tasks. Building upon insights from Ding et al. (2022); Houlsby et al. (2019), they proposed incorporating a lightweight, adaptable module, called an adapter, into a pre-trained and fixed encoder. This addition enables the efficient continual learning of new knowledge by adapting the model for each task to grasp a new set of classes. To address the task identification challenge in continual learning scenarios, task-specific heads were introduced. These heads include an extra class that assimilates all previously learned old classes. Another classification study (Chee et al., 2023) introduced a dynamic architecture featuring expanding representations. Their approach involved incorporating a high-level feature extractor for each task while maintaining a shared low-level feature extractor. This architecture aims to retain common features while learning task-specific characteristics incrementally, asserting the improved preservation of information during each step. In addition, another work by Mousser et al. (2022) proposed an Incremental Deep Tree (IDT) framework tailored for biological image classification. The IDT framework adopted a hierarchical tree-like structure, constructing a new model for each new class by appending branches linked to previously learned tasks. These branches serve to regulate previously acquired knowledge by leveraging replay data and mitigating the issue of catastrophic forgetting. As a result, the framework continuously updates knowledge, enabling the root model to predict new classes while retaining past information and ensuring consistent accuracies for each learned class. Xie et al. (2023) proposes a simple yet effective approach by training batch normalization layer for every task while keeping the rest frozen.

### v2

@@ III-A2 Functional Regularization
[48] proposed learning an inverse function that maps the model’s logits to the feature space and uses it to regularize the feature space of the model in the next task. They demonstrated the effectiveness of their method in a multidomain, single-task X-ray classification scenario. An interesting domain incremental study provided a dual-distillation method combined with a replay to mitigate forgetting [49]. Dual-distillation is carried out using the logits from the old and a new fully fine-tuned model to train an updated model that performs well on new and old tasks. Recent work on manifold learning [50] used a reformer network to augment any input domain image to lie in the same data manifold. This, combined with various regularization-based methods, was shown to be effective in reducing forgetting due to domain shifts in the clinical data. Moreover, a multimodal study by [51] on ultrasound and eye-tracking data utilized distillation to train a model on a continual stream of data, whereas [51] proposed an architecture that leverages the simultaneous availability of two or more datasets to learn a disentanglement between the content and domain in an adversarial manner. This allows a domain-invariant content representation that lays the basis for the continual semantic segmentation of MRI images.[52] use a probability-map alignment scheme to integrate the previous and current map for a segmentation task. They further used knowledge distillation to transfer the knowledge from the integrated probability map to the current model, thereby improving the segmentation performance of the current model. [53] presents a semi-supervised continual learning method combined with a contrastive loss. Their approach was to first train an auto-encoder with a small amount of labeled data in the usual supervised manner, followed by an unsupervised adaptive regularization-based contrastive learning approach.
@@
Another work by [54] introduced a method for multi-organ segmentation in CT images. Their methodology uses lightweight memory modules and a tailored regularization loss function, allowing the model to effectively utilize replayed examples to prevent forgetting. Here, the introduced loss function constrains the learned feature space to maintain the feature of an old class close to its mean representation within the memory module, while ensuring that the features of a new class are distinct from all representations of the old class. Furthermore, [55] proposed a multi-scale multitask CL framework. Their method combined contrastive learning with distillation-based regularization and data rehearsal to enhance learning without catastrophic forgetting. [56] used feature whitening to regularize the important domain in-variant features in a domain incremental cardiac image segmentation study. This feature regularization was augmented by a generative feature replay, where they proposed a base conditional generator along with a style module to produce images for replay. Furthermore, [57] introduced the use of a mixed-curvature embedding space for knowledge distillation to prevent catastrophic forgetting. The authors employed a combination of Euclidean and hyperbolic spaces to project the learned feature embeddings, emphasizing their importance for complex data, especially in medical imaging. Another recent study by [58] used a Kullback–Leibler (KL) divergence-based loss function to preserve the models’ knowledge by penalizing the change in output of the new and old models. More recently, the Fourier test–time adaptation (FTTA) method was presented for classifying medical images with domain gaps [59]. This technique utilizes a Fourier-based domain adjustment strategy along with a multilevel consistency approach to regularize the model’s output. The proposed method is based on a multilevel consistency measurement designed for self-correcting predictions.
Overall, regularization-based methods offer an efficient and privately protected approach to addressing catastrophic forgetting in CL. These methods do not introduce significant computational overhead or privacy concerns, making them initially well-suited for real-life medical applications. However, regularization-based methods also have notable limitations. Empirical evidence shows that these methods often lead to significant performance deterioration over time, particularly when learning diverse and distinct tasks. The trade-off between plasticity and stability can result in models that become increasingly rigid or too plastic, ultimately limiting their practical use. Additionally, as more tasks are learned, the gap between the performance of regularization-based models and the ideal of maintaining individual models grows wider. These challenges make regularization-based methods less effective for long-term use in dynamic environments. Recent efforts to combine regularization-based approaches with other CL techniques have shown promise in mitigating these issues, which could be a key direction for making regularization-based CL methods more practical in medical applications.
III-B Replay-Based Approaches
@@
[60] use Memory-augmented neural networks(MANN) for continuous domain adaptation for the task of semantic segmentation. They proposed a single framework for domain adaptation in which images similar to the target image and their respective textual and shape features are fetched using wavelet features of the target image as a query. These context features, along with the target image features, were embedded into the model trained using only the source domain. [61] proposed a single approach for the different variants of lifelong learning, such as multi-task setting and single incremental task setting, aiming to ensure data privacy and mitigate catastrophic forgetting. They employed a pseudo-rehearsal strategy using a finite memory module and ensured class separation using a composite loss. Furthermore, they introduce feature transformers to learn new representations from the old data as well as the new data. [62] proposed a rehearsal method that infers the data shifts by adopting a dynamic memory module. Their memory update strategy was designed based on a fixed set of rules, and the features were replaced with a high-level metric calculated using gram matrices. [63] employs the same strategy but also introduces an additional module called the pseudo-domain (PD) model, which detects clusters of similar styles from the continuous stream of data. These pseudo-domains are seen as proxies for the unknown, real domains, and are used to balance the memory and training process. [64] also proposed another domain incremental learning that leverages vector quantization approach under limited memory constraints to effectively store and replay hidden representations where no clear domain-shift boundaries exist and no a-priori knowledge is possible. [65] introduced a method that integrates medical expertise into disease diagnosis by considering context and medical entity characteristics in order to transfer knowledge to new stages. Their work was among the pioneering studies exploring CL within medical domains, specifically focusing on diagnosing diseases based on clinical notes. Additionally, they introduced a new technique called embedding episodic memory and consolidation (E2MC) to avoid catastrophic forgetting in disease diagnosis tasks. [66] delve into continual learning within dermatology, employing four distinct methods: naive, elastic weight consolidation, averaged gradient episodic memory, and experience replay. These methodologies allow adaptation to new conditions while safeguarding previously acquired knowledge in the field.
@@
To overcome the problem of insufficient sensitive patient data for digital biomarker model training owing to privacy laws, [67] proposed a Privacy-Preserving Synthetic Data Generation pipeline that uses a GAN to generate anonymous medical samples. Privacy is ensured using a Siamese Neural Network (SNN) with triplet loss to evaluate the generated data. The effectiveness of this method was demonstrated using a Respiratory Sound Classifier that collects respiratory symptom information and is relevant for conditions such as Asthma, COPD, or COVID-19.
@@
Similarly, [68] pioneered a generative replay-based approach in semantic segmentation tasks for medical imaging by introducing SegGAN to prevent catastrophic forgetting. This GAN generates both previous images and their corresponding pixel-level labels, addressing privacy limitations while preserving the acquired knowledge for the subsequent learning stages. Additionally, they proposed a unique filtering mechanism that ensures the selection of high-quality generated data by maintaining consistency between the pseudo-labeling and generative replay techniques. [69] propose the use of multiple generators to improve the quality of replayed samples. Instead of using generators to generate higher-resolution images, this study uses a fixed feature extractor and focuses on feature generation, which requires fewer parameters and fewer computations. [70] introduced a fresh approach for CL in gesture recognition, specifically for adding new gesture classes. This method focuses on managing out-of-distribution (OOD) memory and improving rehearsal to consistently learn new gestures. Additionally, it incorporates an energy-based loss to penalize out-of-distribution gesture examples and further refine the model during training. [71] introduces an innovative data-free framework for incremental class learning. This approach initially creates synthetic data from the model trained on past classes to form a class impression, and then updates the model by merging the synthesized data with new class data. They employ various loss functions, such as cosine normalized cross-entropy, margin loss, and intra-domain contrastive loss, to tackle class imbalances, differentiate between previous and new classes, and ensure the model’s generalization from synthetic to real data. [72] introduce a new method, SMG (Synchronous Memorizability and Generalizability) -learning, to enhance both memorization and generalizability. It presents the Synchronous Gradient Alignment (SGA) objective of improving network memorization and adaptability across previous and new sites. To streamline optimization without excessive computational load, the dual-meta algorithm was developed. Moreover, to enhance rehearsal effectiveness, the replay buffer is modified to include diverse data from various sites, thus minimizing redundancy.
@@
[73] introduce a novel solution to improve image selection to create the memory bank by ranking and selecting images based on their contribution to the learning process. In situations such as medical applications where storing previous data is limited by privacy regulations, [74] presented Generative Appearance Replay for continual Domain Adaptation (GarDA), an approach employing generative replay techniques. GarDA facilitates the gradual adjustment of a segmentation model to new domains by utilizing unlabeled data. Similarly, [75] proposed a generative feature-driven image replay with a dual-purpose discriminator to align features for continual learning in unlabeled domains, achieving state-of-the-art results in histopathology datasets. Unlike single-step unsupervised domain adaptation (UDA), continual adaptation across various domains enables the utilization and consolidation of knowledge from different domains throughout consecutive learning phases. [76] proposes a data selection strategy using Variational autoencoders (VAE) and adversial network. Through this approach, they retain training efficiency by always selecting a fixed subset size from the entire dataset. [77] proposes a continual learning for multicenter studies, eliminating the need for a central server and mitigating the risk of catastrophic forgetting of previously acquired knowledge. The framework involves a method selection process for continual learning, acknowledging that no single method suits all datasets and utilizes fake data generated by a generative adversarial network to evaluate the methods.
Replay-based techniques can be applied in medical settings; however, they bring additional challenges related to model interpretability, regulatory validation, and ethical concerns, all of which must be addressed to ensure the safety and effectiveness of AI-driven healthcare solutions. The primary issue with rehearsal-based methods is the consideration of data privacy, as storing subsets of data may not be feasible due to privacy regulations. This challenge can be mitigated through pseudo-rehearsal or generative rehearsal techniques, where synthetic samples are generated instead of storing actual data. Overcoming these challenges is essential for fully leveraging the potential of continual learning techniques in healthcare while safeguarding patient privacy and safety.
III-C Dynamic Model Approaches
@@
Previous methods were either limited by the model’s parameter change in subsequent tasks or relied on rehearsal techniques. A more recent and promising solution involves intelligent introduction of new parameters for each task (Figure 7). In other words, shared features are maintained among tasks to avoid repeated training, similar to ensembling. Moreover, employing a set of new parameters enables more effective learning of task-specific features.
@@
Dynamic models offer a direct solution by allocating specific segments of the model to distinct tasks. A study by [78] addressed domain shifts in lifelong learning scenarios using domain-specific batch normalization layers for MRI segmentation. Variance in MRI image intensities, caused by multiple factors, leads to this domain shift [79, 80].
@@
However, another challenge arises when we lack knowledge about the task or which sub-network to choose during inference for particular data. For example, if a model is trained on images from three different scanners, determining the domain during inference requires image metadata to activate the corresponding model segment. [81] tackled this issue by training separate UNet models for domain segmentation. To discern the domain during inference, they utilized separate Variational Autoencoders (VAEs) for each domain and selected the domain with the minimal VAE reconstruction loss. Another study by [82] introduces a task-agnostic continual learning approach applicable across diverse deep learning architectures. Their method involved utilizing the Mahalanobis distance to identify domain changes. Upon detecting such changes, they expand the network, ensuring a nonlinear expansion. During inference, the network leverages this distance measurement to determine task ID, thereby facilitating efficient task identification. Similarly, [83] selected a network based on a distance-based metric for each test dataset. In contrast, [84] used different prompts for each task and used a matching strategy to choose the correct prompt at inference.
@@
Recently, advanced techniques have been applied to address the challenges of multiorgan segmentation ([85]). These methods operate in a class incremental setting, requiring the model to segment all the regions observed in previous tasks during inference. [86] highlights the significance of dynamically expanding models to encompass new data and classes for segmenting multiple organs and tumors. Their architecture features lightweight class-specific heads that replace traditional output layers, enabling separate predictions for both new and previously learned classes. Leveraging extensive image-text co-training, the authors embedded contrastive language-image pretraining (CLIP) embeddings into organ-specific heads to capture each class’s semantic essence. Research by [87] introduced an innovative incremental transfer learning framework designed to progressively address multi-site medical image segmentation tasks. They proposed strategies at both the model and site levels aimed at incremental training to enhance segmentation, improve generalization, and boost transfer performance, particularly in settings constrained by limited clinical resources.
@@
[88] introduced an innovative architectural framework for semantic segmentation. Their approach learns a unified deep segmentation model capable of segmenting 143 whole-body organs by training a single unified deep segmentation model. For each segmentation task, a trained encoder remained fixed, whereas a new decoder was gradually added. To handle model complexity, they utilized a progressive pruning strategy involving neural architecture search and teacher-student knowledge distillation. This strategy helps maintain performance while controlling the size of the model.
@@
The practice of dynamically adjusting models extends to classification tasks, as observed in a recent study by [89] on various medical imaging classification tasks. Building upon insights from [90, 91], they proposed incorporating a lightweight, adaptable module, called an adapter, into a pre-trained and fixed encoder. This addition enables the efficient continual learning of new knowledge by adapting the model for each task to grasp a new set of classes. To address the task identification challenge in continual learning scenarios, task-specific heads were introduced. These heads include an extra class that assimilates all previously learned old classes. Another classification study [92] introduced a dynamic architecture featuring expanding representations. Their approach involved incorporating a high-level feature extractor for each task while maintaining a shared low-level feature extractor. This architecture aims to retain common features while learning task-specific characteristics incrementally, asserting the improved preservation of information during each step. In addition, another work by [93] proposed an Incremental Deep Tree (IDT) framework tailored for biological image classification. The IDT framework adopted a hierarchical tree-like structure, constructing a new model for each new class by appending branches linked to previously learned tasks. These branches serve to regulate previously acquired knowledge by leveraging replay data and mitigating the issue of catastrophic forgetting. As a result, the framework continuously updates knowledge, enabling the root model to predict new classes while retaining past information and ensuring consistent accuracies for each learned class. [94] proposes a simple yet effective approach by training batch normalization layer for every task while keeping the rest frozen.